OpenStudy (anonymous):

graph each function by first writing it in vertex form. then give maximum or minimum of the function and identify its zeros. 1.) f(x)=x^2-6x+9 2.)f(x)=x^2-2x-3 3.f(x)=-7x^2-14x 4.)f(x)=3x^2-12x+9

OpenStudy (triciaal):

the equation is already in general/standard form Ax^2 + By +C = 0 vertex form a(x-h)^2 + k = 0 the x coordinate is -b/2a let this be h then find 2c -(b/2a)^2 let this be k

OpenStudy (triciaal):

the maximun or minimum occurs when f(x) = 0 to identify zeros x = h and k = 0

OpenStudy (triciaal):

1.) f(x)=x^2-6x+9 a =1, b = -6 and c= 9 -b/2a = -(-6)/2*1 = 3 f(3) = 3^2 -6*3 + 9 =0 f(x) = (x-3)^2 zero when x = 3

OpenStudy (triciaal):

2.)f(x)=x^2-2x-3 a = 1 b = -2 c = -3 -b/2a = - - 2/2*1 = 1 k =-3 - 1^2 = -3 -1 =-4 f(x) =(x-1)^2 -4 when f(x) = 0 then (x-1)^2 =4 x -1 = +/- 2 x = 2 + 1 , x = -2 + 1 x = 3, -1

OpenStudy (triciaal):

the y -coordinate of a parabola is f(-b/2a)

OpenStudy (triciaal):

try the rest